# Federated Instruction Tuning of LLMs with Domain Coverage Augmentation

### v1

@@ 6 Conclusion
We reveal that for FedDIT, the model performance is not linearly correlated with data heterogeneity, but is significantly impacted by cross-client domain coverage. To address this, we propose FedDCA, a domain-specific instruction augmentation method that maximizes domain coverage through greedy client center selection and retrieval-based augmentation. Additionally, FedDCA uses a heterogeneous encoder to reduce client-side computation overhead. Experiments across four domains demonstrate the effectiveness of FedDCA and the efficiency of FedDCA. Further analysis shows that as fine-tuning progresses, the risk of privacy leakage from augmented data decreases or converges.
@@ A.1 Related Work
Federated Instruction Tuning. Instruction tuning has been widely applied across various application areas of large language models (LLM), serving as a key technique to enhance the capabilities and controllability of LLM (Zhang et al. (2024b); Wei et al. (2022)). Recently, federated instruction tuning (FedIT) has emerged as an effective strategy for the distributed optimization of LLMs, leveraging federated learning (FL) protocols to improve the handling of privacy-sensitive tasks in real-world scenarios. So far, several FedIT frameworks(Ye et al. (2024b; a)) have been established to evaluate the effectiveness of FedIT across multiple datasets, tasks, and FL methods. While these platforms provide a foundation for research, they have not yet introduced more complex federated algorithms and deeply investigate the challenging problems and factors affecting FedIT, which are crucial for advancing this field.
@@
PrivateLoRA(Wang et al. (2023b)) addresses privacy and efficiency issues by exploiting the low-rank properties of residual activations to reduce communication costs, significantly lowering the communication overhead through collaborative computation between the server and clients while effectively maintaining the privacy of local data. While FewFedPIT(Zhang et al. (2024d)) focuses on the few-shot learning setting in FedIT, using self-generated data by pre-trained LLMs locally to mitigate the paucity of data and first discussing memory extraction attacks within FedIT.
@@
Domain Instruction Augmentation. In the real world, there is an urgent need for training LLMs with specific functionalities (e.g., reasoning capabilities) or domain-specific LLMs (e.g., code(Nijkamp et al. (2023); Luo et al. (2023b)), medical(2024-Zhang-AlpaCare-), financial(Yang et al. (2023b; a); Zhang et al. (2023a); Wu et al. (2023)), mathematical(Yue et al. (2024); Luo et al. (2023a))). Existing works tend to use open-source domain-specific instruction tuning datasets for training. However, the target domain may not always have corresponding ready-made domain-specific instruction datasets. Even if they exist, these datasets are often limited in scale.
@@
Several studies investigate domain-specific instruction augmentation, which can be categorized into three aspects: 1) Reusing human-curated public datasets(Wang et al. (2023a); Zhang et al. (2023c); Jiao et al. (2023); Lee et al. (2024); Xia et al. (2024)). For instance, Parrot(Jiao et al. (2023)) enhances translation capabilities of LLMs in chat by converting bilingual sentences into instruction-following datasets. Furthermore, works like INSTA(Lee et al. (2024)) and LESS(Xia et al. (2024)) attempt efficient domain-specific instruction augmentation via dense retrieval. INSTA(Lee et al. (2024)) uses instructions without responses for effective retrieval. LESS(Xia et al. (2024)) assumes access to a validation set and uses the warmup LLM’s gradients of the train and validation sets for retrieval. 2) Using seed tasks for self-instruct(Luo et al. (2023b); Wan et al. (2023)): Explore-Instruct(Wan et al. (2023)), for example, employs activate exploration to tree-model domain tasks from both depth and breadth, increasing seed instructions’ coverage of the domain, and subsequently generating broader in-domain instructions through self-instruct. 3) Scaling instructions from the web: Recent works(Yue et al. (2024); Zhou et al. (2024)) highlight the immense potential of mining naturally occurring instructions from the internet. Compared to generated data, web-mined instructions exhibit less bias and greater diversity. MAmmoTH2(Yue et al. (2024)) firstly retrieves domain-relevant texts and employs a LLM to extract Q-A pairs, further refining them into instruction-response pairs. Jiuzhang3.0(Zhou et al. (2024)) distills GPT’s instruction generation capabilities into a smaller model and then uses it to generate instructions from the internet. In conclusion, models fine-tuned with augmented instructions have shown promising domain-specific capabilities.
Inspired by INSTA(Lee et al. (2024)), this paper uses a mixed dataset composed of code, medical, financial, mathematical, and general datasets to represent public data and does domain-specific instruction augmentation based on the client’s local instructions.
@@ A.2 What Truly Counts in FedDIT
This section will first analyze the correlation between different non-iid levels and the model’s performance with separate experiments for both single and multiple domains in Section A.2.1. Further, to demonstrate the impact of non-iid on FedDIT with different test sets, we compare the performance of the global model trained on augmented data based on iid and non-iid cross-client data distribution. Additionally, we suggest that domain coverage is a key factor for FedDIT in Section A.2.2.
@@ A.2.1 Data Heterogeneity Is Not Matter In FedDIT
Following the traditional approach of constructing different degrees of non-iid which are widely used in federated learning(Wang et al. (2020); Yurochkin et al. (2019)), we adopt Dirichlet distribution to construct various heterogeneity and use k-means with the cluster num $\xi=100$ to pseudo labeling instructions. Dirichlet distribution is affected by the hyperparameter $\alpha$, which enhances heterogeneity with a smaller $\alpha$ and decreases heterogeneity with a larger $\alpha$. We choose four widely used heterogeneity, which are $\alpha=[0.01,0.1,1,10]$(Ye et al. (2023); Zhang et al. (2023b); li2021model). Figure LABEL:fig:data_heterogeneity shows a visualization of the data distribution with different heterogeneity on the code dataset.
We perform instruction tuning both on single-domain and multi-domain with different heterogeneity. Specifically, for the single domain, we only use the in-domain data and then perform FedIT, and for multi-domain, we use the mixed dataset which consists of four domain-specific instruction datasets and a general instruction dataset, which are described in Appendix A.5. The training details are shown in Appendix A.6. Results for single domain and multi-domain are shown in Table 6. Both single-domain and multi-domain clearly show that the performance of LLM does not decrease due to the increase of data heterogeneity, but shows a nonlinear correlation, which indicates that the performance of LLM does not directly depend on data homogeneity, and exists other factors that play a key role.
@@ A.2.2 Domain Coverage: A Key Factor in FedDIT
Explore-Instruct(Wan et al. (2023)) enhances the coverage of domain-specific seed tasks through active exploration, then uses the self-instruct method for instruction data augmentation. This approach highlights the impact of domain coverage on domain-specific instruction tuning. Inspired by Explore-Instruct, we attempt to conduct more in-depth and extensive experiments to study the effect of domain coverage on FedDIT. Firstly we define the domain coverage of cross-client data in the FL setting. Assume the dataset of in-domain data $D^{d}$ represents the latent data distribution of this domain and the cross-client data is defined as $D^{c}=\cup_{k=1}^{N}\left(D_{k}^{l}\cup D_{k}^{g}\right)$. Inspired by the facility location function(Cornuejols et al. (1983)), we define the domain coverage of $D^{c}$ respect to $D^{d}$ as follows:
@@
where $sim(d,v)$ is the similarity between $d$ and $v$. Specifically, we use the cosine similarity function in FedDCA. Note that we only use the in-domain data in $D^{c}$ to calculate the domain coverage, because the out-of-domain data would mislead the domain coverage evaluation, as shown in Figure 7.
@@ A.3 Direct domain data retrieval
In this section, we first show the detail of instruction-based dense retrieval in Appendix A.3.1, which is both used in direct retrieval and FedDCA. Then we explain the direct retrieval algorithm in Appendix A.3.2.
@@ A.3.2 Direct Retrieval
The direct domain data retrieval only utilizes instructions of the client’s local data without responses to perform the retrieval-based domain data augmentation. The detailed algorithm is described in Algorithm 2. In specific, for each client, we first encode the local instructions by the encoder model $w_{enc}$. Then we use the k-means algorithm to cluster the embeddings into $\xi$ clusters. Then the $\xi$ cluster centers are sent to the server and do the retrieval-based domain data augmentation. Then the retrieved public data is sent to the client for instruction tuning.
@@ A.4 Projector
To reduce the computational overhead of clients, we propose a heterogeneous encoder method called FedDCA. Inspired by LLaRA(Liao et al. (2023)), we utilize a projector to do the dimension alignment between the small encoder on the client side and the large decoder on the server side. As shown in Figure 8, where $d_{l}$ is the dimension of the client-side encoder, $d_{g}$ is the dimension of the server-side encoder.
@@ A.5 Train and Test Dataset Information
Here we provide the train and test dataset details of code (CodeAlpaca [^2] https://huggingface.co/datasets/sahil2801/CodeAlpaca-20k), medical (MedAlpaca [^3] https://huggingface.co/datasets/medalpaca), financial (FinGPT(Yang et al. (2023a))) and mathematical (MathInstruct [^4] https://huggingface.co/datasets/TIGER-Lab/MathInstruct) domain, respectively.
@@
As shown in Table 8, the public data consists of four domain-specific instruction datasets and a general instruction dataset, which are CodeAlpaca, MedAlpaca, FinGPT, MathInstruct, and Alpaca, respectively. For each domain’s FedDIT, we randomly select 1000 samples from the in-domain instruction and split them into 10 shards as each client’s local dataset. The rest instructions are used as the public dataset $D^{p}$.
@@
Table 9 shows the detail of the test set used for evaluation, HumanEval is used for the code domain, MMLU-Med is used for the medical domain, FPB, FiQA and TFNS are used for the financial domain, and GSM8K is used for the mathematical domain. Specifically, MMLU-Med uses subjects anatomy, clinical_knowledge, college_biology, college_medicine, medical_genetics and professional_medicine in MMLU.
@@ A.6 Implementation Details
We consider FedDIT in the cross-device scenario, $N$ = 10 clients, $\mathcal{R}$ = 30 rounds, where we randomly sample 2 clients to be available for each round. Then each available client performs FedDIT for 10 steps with AdamW optimizer and the batch size $B=32$ in a round. The initial learning rate is $55e-$ 5 with a cosine learning rate scheduler. Our experiment utilizes the widely used LLM, Llama3-8B [^5] https://huggingface.co/meta-llama/Meta-Llama-3-8B as the base model with 2048 max sequence length and adopts LoRA tuning method. The rank of LoRA is 16 and the scalar alpha is 16. For k-means(Vardakas & Likas (2023)), we set cluster num $\xi$ = 10 and for FedDCA we set the similarity threshold $\alpha$ = 0.7. For FedDCA, we set the temperature parameter $\tau$ = 0.5 for contrastive learning(Khosla et al. (2020)). We utilize bge-large-en-v1.5666https://huggingface.co/BAAI/bge-large-en-v1.5 as both the client and server’s encoder as default, which outputs embeddings of 1024 dimensions. While we utilize all-MiniLM-L6-v2777https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 as the client’s small encoder which outputs embeddings of 384 dimensions and bge-large-en-v1.5 as the server’s encoder for FedDCA.
@@ A.7 Prompts used in the Self-Instruct data generation
To generate the Self-Instruct data, we prompt GPT-3.5 to generate the instruction with the designed prompt in Figure 9. Specifically, we randomly sample two examples from the client’s local data to guide GPT-3.5 generating the in-domain instruction, and sample one example from the client’s local data for one-shot in-context learning to guide GPT-3.5 generating responses into the example’s format.
@@ A.9 Further Analysis
To further study the effect of different hyperparameters of FedDCA, we undertake a thorough analysis including various retrieval amounts, and different k-means cluster numbers $\xi$. In addition, we perform the ablation study on whether using the similarity threshold $\alpha$ in the greedy client center selection.
@@
Figure 11: Effect of different retrieval amounts on the performance of FedDCA and its domain coverage. We show the results on four domains separately. Here we use the FPB test set to evaluate the performance in the financial domain.
@@
Ablation Study. We conduct the experiment with FedDCA w/o similarity threshold $\alpha$ on the four domains based on the FedAvg FL strategy. The performance and the corresponding domain coverage are shown in Table 10, where FedDCA without the similarity threshold $\alpha$ is marked as FedDCA. The result shows that the performance of FedDCA with similarity threshold $\alpha$ is slightly better than FedDCA without using $\alpha$ in code, financial, and mathematical domains, as the similarity scores in the medical domain are relatively lower. We show the similarity score distribution of the four domains in Figure LABEL:fig:similarity_score. For each domain, we plot each similarity score’s distribution of 10 clients. The similarity score is computed between the the selected client center and the public data, then we show the similarity score distribution using the histogram plot.

### v3

@@ 6 Conclusion
We reveal that in the context of FedDIT, the model performance is not linearly correlated with data heterogeneity but is significantly impacted by the cross-client domain coverage. In response, we propose a novel FedDIT method called FedDCA, which optimizes the domain coverage through greedy client center selection and retrieval-based instruction augmentation. Additionally, FedDCA leverages heterogeneous encoders to reduce the client-side computation overhead and improve system scalability. Experiments across four domains demonstrate the effectiveness of FedDCA and the efficiency of FedDCA. Further privacy analysis indicates that as fine-tuning advances, the risk of private data leakage diminishes or converges in FedDIT with instruction augmentation.
@@ A.1 Related Work
Federated Instruction Tuning. Instruction tuning has been widely applied across various application areas of large language models (LLM), serving as a key technique to enhance the capabilities and controllability of LLM (Zhang et al., 2024b; Wei et al., 2022). Recently, federated instruction tuning (FedIT) has emerged as an effective strategy for the distributed optimization of LLMs, leveraging federated learning (FL) protocols to improve the handling of privacy-sensitive tasks in real-world scenarios. So far, several FedIT frameworks (Ye et al., 2024b; a) have been established to evaluate the effectiveness of FedIT across multiple datasets, tasks, and FL methods. While these platforms provide a foundation for research, they have not yet introduced more complex federated algorithms and deeply investigate the challenging problems and factors affecting FedIT, which are crucial for advancing this field.
@@
PrivateLoRA (Wang et al., 2023b) addresses privacy and efficiency issues by exploiting the low-rank properties of residual activations to reduce communication costs, significantly lowering the communication overhead through collaborative computation between the server and clients while effectively maintaining the privacy of local data. While FewFedPIT (Zhang et al., 2024d) focuses on the few-shot learning setting in FedIT, using self-generated data by pre-trained LLMs locally to mitigate the paucity of data and first discussing memory extraction attacks within FedIT.
@@
Domain Instruction Augmentation. In the real world, there is an urgent need for training LLMs with specific functionalities (e.g., reasoning capabilities) or domain-specific LLMs (e.g., code (Nijkamp et al., 2023; Luo et al., 2023b), medical (2024-Zhang-AlpaCare-), financial (Yang et al., 2023b; a; Zhang et al., 2023a; Wu et al., 2023), mathematical (Yue et al., 2024; Luo et al., 2023a)). Existing works tend to use open-source domain-specific instruction tuning datasets for training. However, the target domain may not always have corresponding ready-made domain-specific instruction datasets. Even if they exist, these datasets are often limited in scale.
@@
Several studies investigate domain-specific instruction augmentation, which can be categorized into three aspects: 1) Reusing human-curated public datasets (Wang et al., 2023a; Zhang et al., 2023c; Jiao et al., 2023; Lee et al., 2024; Xia et al., 2024). For instance, Parrot (Jiao et al., 2023) enhances translation capabilities of LLMs in chat by converting bilingual sentences into instruction-following datasets. Furthermore, works like INSTA (Lee et al., 2024) and LESS (Xia et al., 2024) attempt efficient domain-specific instruction augmentation via dense retrieval. INSTA (Lee et al., 2024) uses instructions without responses for effective retrieval. LESS (Xia et al., 2024) assumes access to a validation set and uses the warmup LLM’s gradients of the train and validation sets for retrieval. 2) Using seed tasks for self-instruct (Luo et al., 2023b; Wan et al., 2023): Explore-Instruct (Wan et al., 2023), for example, employs activate exploration to tree-model domain tasks from both depth and breadth, increasing seed instructions’ coverage of the domain, and subsequently generating broader in-domain instructions through self-instruct. 3) Scaling instructions from the web: Recent works (Yue et al., 2024; Zhou et al., 2024) highlight the immense potential of mining naturally occurring instructions from the internet. Compared to generated data, web-mined instructions exhibit less bias and greater diversity. MAmmoTH2 (Yue et al., 2024) firstly retrieves domain-relevant texts and employs a LLM to extract Q-A pairs, further refining them into instruction-response pairs. Jiuzhang3.0 (Zhou et al., 2024) distills GPT’s instruction generation capabilities into a smaller model and then uses it to generate instructions from the internet. In conclusion, models fine-tuned with augmented instructions have shown promising domain-specific capabilities.
To obtain a well-performing LLM in the specific domain within the distributed environment, we utilize a multi-domain dataset as public data on the server side and perform domain-specific instruction augmentation based on the client’s local instructions.
@@ A.2 What Truly Counts in FedDIT
This section will first analyze the correlation between different non-iid levels and the model’s performance with separate experiments for both single and multiple domains in Section A.2.1. Further, to demonstrate the impact of non-iid on FedDIT, we compare the performance of the global model trained on augmented data based on iid and non-iid cross-client data distribution. Additionally, we suggest that domain coverage is a key factor for FedDIT in Section A.2.2.
@@ A.2.1 Data Heterogeneity Is Not Matter In FedDIT
Following the traditional approach of constructing different degrees of non-iid, which are widely used in federated learning (Wang et al., 2020; Yurochkin et al., 2019), we adopt Dirichlet distribution to construct various heterogeneity and use k-means with the cluster num $\xi=100$ to pseudo labeling instructions. Dirichlet distribution is affected by the hyperparameter $\alpha$, which enhances heterogeneity with a smaller $\alpha$ and decreases heterogeneity with a larger $\alpha$. We choose four widely used heterogeneity, which are $\alpha=[0.01,0.1,1,10]$ (Ye et al., 2023; Zhang et al., 2023b; li2021model). Figure LABEL:fig:data_heterogeneity shows a visualization of the data distribution with different heterogeneity on the code dataset.
We perform instruction tuning both on single-domain and multi-domain with different heterogeneity. Specifically, for the single domain, we only use the in-domain data and then perform FedIT. For multi-domain, we use the mixed dataset, which consists of four domain-specific instruction datasets and a general instruction dataset, which are described in Appendix A.5. The training details are shown in Appendix A.6. Results for single domain and multi-domain are shown in Table 6. Both single-domain and multi-domain clearly show that the performance of LLM does not decrease due to the increase of data heterogeneity but shows a nonlinear correlation, which indicates that the performance of LLM does not directly depend on data heterogeneity and other factors that play a key role.
@@ A.2.2 Domain Coverage: A Key Factor in FedDIT
Explore-Instruct (Wan et al., 2023) enhances the coverage of domain-specific seed tasks through active exploration, then uses the self-instruct method for instruction data augmentation. This approach highlights the impact of domain coverage on domain-specific instruction tuning. Inspired by Explore-Instruct, we attempt to conduct more in-depth and extensive experiments to study the effect of domain coverage on FedDIT. Firstly, we define the domain coverage of cross-client data in the FL setting. Assume the dataset of in-domain data $D^{d}$ represents the latent data distribution of this domain and the cross-client data is defined as $D^{c}=\cup_{k=1}^{N}\left(D_{k}^{l}\cup D_{k}^{g}\right)$. Inspired by the facility location function (Cornuejols et al., 1983), we define the domain coverage of $D^{c}$ respect to $D^{d}$ as follows:
@@
where $sim(d,v)$ is the similarity between $d$ and $v$. Specifically, we use the cosine similarity function in FedDCA. Note that we only use the in-domain data in $D^{c}$ to calculate the domain coverage because the out-of-domain data would mislead the domain coverage evaluation, as shown in Figure 7.
@@ A.3 Direct domain data retrieval
In this section, we first show the detail of instruction-based dense retrieval in Appendix A.3.1, which is both used in direct retrieval and FedDCA. Then, we explain the direct retrieval algorithm in Appendix A.3.2.
@@ A.3.2 Direct Retrieval
The direct domain data retrieval only utilizes instructions of the client’s local data without responses to perform the retrieval-based domain data augmentation. The detailed algorithm is described in Algorithm 2. For each client, we start by encoding the local instructions using the encoder model $w_{enc}$. Next, we apply the k-means algorithm to cluster the embeddings into $\xi$ clusters. Then $\xi$ cluster centers are sent to the server for retrieval-based domain data augmentation. Subsequently, the retrieved public data is sent to the client for instruction tuning.
@@ A.4 Projector
To reduce the computational overhead of clients, we propose a heterogeneous encoder method called FedDCA. We utilize a projector to perform the dimension alignment between the small encoder on the client side and the large decoder on the server side. As shown in Figure 8, where $d_{l}$ is the dimension of the client-side encoder, $d_{g}$ is the dimension of the server-side encoder.
@@ A.5 Train and Test Dataset Information
Here we provide the train and test dataset details of code (CodeAlpaca [^2] https://huggingface.co/datasets/sahil2801/CodeAlpaca-20k), medical (MedAlpaca [^3] https://huggingface.co/datasets/medalpaca), financial (FinGPT (Yang et al., 2023a)) and mathematical (MathInstruct [^4] https://huggingface.co/datasets/TIGER-Lab/MathInstruct) domain, respectively.
@@
As shown in Table 8, the public data consists of four domain-specific instruction datasets and a general instruction dataset, which are CodeAlpaca, MedAlpaca, FinGPT, MathInstruct, and Alpaca, respectively. For each domain’s FedDIT, we randomly select 1000 samples from the in-domain instruction and split them into 10 shards as each client’s local dataset. The rest of the instructions are used as the public dataset $D^{p}$.
@@
Table 9 shows the detail of the test set used for evaluation: HumanEval is used for the code domain; MMLU-Med is used for the medical domain; FPB, FiQA and TFNS are used for the financial domain, and GSM8K is used for the mathematical domain. Specifically, MMLU-Med uses subjects anatomy, clinical_knowledge, college_biology, college_medicine, medical_genetics and professional_medicine in MMLU.
@@ A.6 Implementation Details
We consider FedDIT in the cross-device scenario, $N$ = 10 clients, $\mathcal{R}$ = 30 rounds, where we randomly sample 2 clients to be available for each round. Then, each available client performs FedDIT for 10 steps with AdamW optimizer, and the batch size is $B=32$ in a round. The initial learning rate is $55e-$ 5 with a cosine learning rate scheduler. Our experiment utilizes the widely used LLM, Llama3-8B [^5] https://huggingface.co/meta-llama/Meta-Llama-3-8B as the base model with 2048 max sequence length and adopts LoRA tuning method. The rank of LoRA is 16, and the scalar alpha is 16. For k-means (Vardakas & Likas, 2023), we set cluster num $\xi$ = 10 and for FedDCA we set the similarity threshold $\alpha$ = 0.7. For FedDCA, we set the temperature parameter $\tau$ = 0.5 for contrastive learning (Khosla et al., 2020). We utilize bge-large-en-v1.5666https://huggingface.co/BAAI/bge-large-en-v1.5 as both the client and server’s encoder as default, which outputs embeddings of 1024 dimensions. While we utilize all-MiniLM-L6-v2777https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 as the client’s small encoder which outputs embeddings of 384 dimensions and bge-large-en-v1.5 as the server’s encoder for FedDCA.
@@ A.7 Prompts used in the Self-Instruct data generation
To generate the Self-Instruct data, we prompt GPT-3.5 to generate the instruction with the designed prompt in Figure 9. Specifically, we randomly sample two examples from the client’s local data to guide GPT-3.5 generating the in-domain instruction and one example from the client’s local data for one-shot in-context learning to guide GPT-3.5 generating responses into the example’s format.
@@ A.9 Further Analysis
To further study the effect of different hyperparameters of FedDCA, we undertake a thorough analysis including various retrieval amounts and different k-means cluster numbers $\xi$. In addition, we perform the ablation study on whether using the similarity threshold $\alpha$ in the greedy client center selection.
@@
Figure 11: Effect of different retrieval amounts on the performance of FedDCA and its domain coverage. We show the results on four domains separately. Here, we use the FPB test set to evaluate the performance in the financial domain.
@@
Ablation Study. We conduct the experiment with FedDCA w/o similarity threshold $\alpha$ on the four domains based on the FedAvg FL strategy. The performance and the corresponding domain coverage are shown in Table 10, where FedDCA without the similarity threshold $\alpha$ is marked as FedDCA. The result shows that the performance of FedDCA with similarity threshold $\alpha$ is slightly better than FedDCA without using $\alpha$ in code, financial, and mathematical domains, as the similarity scores in the medical domain are relatively lower. We show the similarity score distribution of the four domains in Figure LABEL:fig:similarity_score. For each domain, we plot each similarity score’s distribution of 10 clients. The similarity score is computed between the selected client center and the public data. Then, we show the similarity score distribution using the histogram plot.